OpenStudy (anonymous):

which equation is NOT an identity? A)tan(-x)=-tan(x) B)tan t(csc t+cot t)= sec t+1 C)sin tcot t=cos t D)cos^2t-sin^2t=-1

OpenStudy (anonymous):

the first one is, since tangent is odd

OpenStudy (anonymous):

does the last one say \[\cos^2(t)-\sin^2(t)=-1\]?

OpenStudy (anonymous):

if so, pick that one since that is certainly not always the case

OpenStudy (anonymous):

yes, it does

OpenStudy (anonymous):

you were right. ty so much

OpenStudy (anonymous):

for example if cos(t)=1 and sin(t) = 0, then 1-0=1 not -1 that is not an identity yw